A grinding wheel, initially at rest, is rotated for 8 seconds with a constant angular acceleration of 5 rad/s^2. The wheel is then brought to rest, with a uniform negative angular acceleration, in 10 revolutions. (a) Determine the angular acceleration needed. (b) How lng did it take to bring it to rest? *


1
Expert's answer
2021-05-13T18:08:27-0400

The velocity of the wheel after the 8 seconds will be

"\omega = \beta_1 \cdot t_1 = 5\,\mathrm{rad/s^2}\cdot 8\,\mathrm{s} = 40\,\mathrm{rad/s}."

10 revolutions correspond to "L=10\cdot 2\pi = 20\pi\,\mathrm{rad}."

This length is also equal to "L = \omega t_2 - \dfrac{\beta_2 t_2^2}{2}" , where "\beta_2" is the acceleration in question (here we write the conditions for modulus of acceleration and put minus sign).

Also "\omega - \beta_2 t_2" should be equal to 0, because the wheel should stop. So we obtain two equations with two unknown values "(\beta_2 \,\mathrm{and}\, t_2)" :

"20\pi = 40 t_2 - \dfrac{\beta_2 t_2^2}{2},\; \; 40 - \beta_2 t_2 = 0."

We express "\beta_2" from the second equation: "\beta_2 = \dfrac{40}{t_2}" and substitute it into the first equation:

"20\pi = 40 t_2 - \dfrac{40 t_2}{2} , \\
20\pi = 20t_2, \\
t_2 = \pi\approx 3.14\,\mathrm{s} \; \Rightarrow \; \beta_2 = \dfrac{40}{\pi}\approx 12.7\,\mathrm{rad/s^2}."

The deceleration is "-12.7\,\mathrm{rad/s^2}."